OpenStudy (anonymous):

n average, a commercial airplane burns about 4.0 × 103 mL of gasoline per second. There can be as many as 5 × 103 airplanes in the air in the United States at any given time. About how many milliliters of gasoline are burned per minute by commercial airplanes in flight in the United States? 1.2 × 10 ______ mL

OpenStudy (mathmate):

Hint: 103 = 100+3 10^3 = \(10^3\) They don't mean the same thing.

OpenStudy (anonymous):

@cc_peach

OpenStudy (anonymous):

i don't know this one... srry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

@pooja195

OpenStudy (mathmate):

Each airplane burns 4*10^3 mL /s = 24*10^3 mL/min. There are 5*10^3 airplanes which burn a total of 5*10^3 * 24*10^3 mL/min =120*10^(3+3) mL /min =120*10^*6 mL/min =1.2*10^? mL/min.